OpenStudy (anonymous):

Quick question on simplifying complex fractions, 1+1/x all over 1-1/x^2

OpenStudy (anonymous):

@kc_kennylau

OpenStudy (anonymous):

\[\frac{ 1+\frac{ 1 }{ x} }{ 1-\frac{ 1 }{ x^2}}\]

OpenStudy (solomonzelman):

I am still disconnected, sorry...

OpenStudy (anonymous):

what do you mean?

OpenStudy (anonymous):

Alright well thank anyway i have to go take my college algebra final :P

OpenStudy (solomonzelman):

1) find the common denominators in the bottom and top fractions. 2) you will get (in terms of variables) \[~~~ \frac{a}{b}~~~\over~~~ \frac{x}{y}~~~\]which is same as \[\frac{a}{b} \div \frac{x}{y}~~~~~->~~~~~~~~~\frac{a}{b} \times \frac{y}{x}~~~~~->~~~~\frac{a \times y}{b \times x}\] 3) see? solve your's.